TigerSharkie:

Which fraction is equivalent to 2 over negative 6 ? (5 points) negative 2 over 6 2 over 6 negative 2 over negative 6 6 over 2

snowflake0531:

Which fraction is equivalent to \(\frac{2}{-6}\) a. \(-\frac{2}{6}\) b. \(\frac{2}{6}\) c. \(\frac{-2}{-6}\) d. \(\frac{6}{2}\) One thing that you should know is that where the negative sign is doesn't matter \[\frac{2}{-6} = \frac{-2}{-6} = -\frac{2}{6}\] By now, you should have your answer